# Fast Crystal Tensor Property Prediction: A General O(3)-Equivariant Framework Based on Polar Decomposition

### v2

@@ 5.1 Experimental setup
Datasets. In this work, we evaluate the performance of GoeCTP on two key tensor property prediction tasks: the second-order dielectric tensor and the fourth-order elastic tensor, respectively. The dataset for dielectric tensor is derived from the data processed by Yan et al. (2024b), sourced from the JARVIS-DFT database. Similarly, the dataset for elastic tensor is obtained from the dft_3d data within the JARVIS-DFT database111 https://pages.nist.gov/jarvis/databases/. The statistical details of the datasets are presented in Table 1. Additionally, in the dielectric tensor dataset, the dielectric tensor is a $3\times 3$ symmetric matrix. Therefore, during prediction, we predict 6 elements of the matrix (see Appendix A.2 for details) and then reconstruct the entire $3\times 3$ symmetric matrix. In the elastic tensor dataset, the elastic tensor is represented using Voigt notation as a $6\times 6$ symmetric matrix, rather than a $3\times 3\times 3\times 3$ fourth-order tensor (Itin & Hehl, 2013; Wen et al., 2024). Therefore, we predict the $6\times 6$ matrix during usage (see Appendix A.2 for details).
Dataset Sample size Fnorm Mean Fnorm STD Unit
Dielectric 4713 14.7 18.2 Unitless
Elastic 25110 306.4 238.4 GPa
Table 1: Dataset statistics.
Baseline Methods. we selected several state-of-the-art methods in the field of crystal tensor property prediction, i.e. MEGNET (Chen et al., 2019; Morita et al., 2020), EGTNN (Zhong et al., 2023), and GMTNet (Yan et al., 2024b), as baseline methods.
@@
Experimental Settings. The experiments were performed using an NVIDIA GeForce RTX 3090 GPU. For benchmarking, we directly utilized the codebases for MEGNET, ETGNN, and GMTNet as provided by Yan et al. (2024b). For each property, the dataset is split into training, validation, and test sets in an 8:1:1 ratio. During model training, Huber loss, AdamW (Loshchilov & Hutter, 2018), a weight decay of $10^{-5}$, and polynomial learning rate decay were employed.
Since we observed some randomness in the results during the training process of the repeated dielectric dataset experiments, we conducted 5 repeated experiments for each method and used the average of the 5 experiment metrics as the final experimental results. The results of the 5 repeated experiments can be found in the Appendix 11.
@@ 5.2 Experimental Results
Predicting Dielectric Tensors. The performance of various models in predicting the dielectric tensor is summarized in Table 2. While GoeCTP and GMTNet show close results in terms of Fnorm and EwT 25%, GoeCTP demonstrates higher values for EwT 10% and EwT 5%, with an improvement of 5% in EwT 5% compared to GMTNet. This indicates that GoeCTP delivers higher-quality predictions compared to GMTNet. Furthermore, as GoeCTP is designed to be a flexible framework, we also evaluated its combination with other models, such as iComFormer (Yan et al., 2024a) and CrystalFormer (Taniai et al., 2024). Detailed results of these combinations are provided in Appendix 11.
@@
Predicting Elastic Tensors. Experimental results on the elastic tensor dataset are presented in Table 3. GoeCTP demonstrated outstanding performance in predicting higher-order, complex tensors, outperforming all baseline methods across every evaluation metric. Notably, it achieved the lowest Fnorm of 107.11 and improved all EwT metrics by an average of 6% compared to GMTNet, emphasizing its general applicability in predicting diverse tensor properties in materials science. Further details on the results of GoeCTP’s combination with other models can be found in Appendix 11.
@@
Verifying the $O(3)$ Equivariance. To evaluate the effectiveness of GoeCTP, we conducted experiments to verify the $O(3)$ equivariance of tensor properties. Specifically, after training GoeCTP, we extracted its Prediction module (i.e., eComFormer) for comparative testing on two different test sets (original test set and augmented test set). All crystals in the original test set were adjusted to the standardized crystal position, while the augmented test set was generated by applying arbitrary $O(3)$ group transformations to all crystals in the original test set. The method for generating the corresponding orthogonal matrices is from Heiberger (1978). We then evaluated both the Prediction module (eComFormer) and GoeCTP on these two datasets and compared the performance metrics.
The results on the dielectric tensor dataset are shown in Table 4. GoeCTP performed equally well on the augmented test set as on the original test set, indicating that it maintains strong $O(3)$ equivariance for tensor properties. In contrast, the performance of the eComFormer method significantly declined on the augmented test set, with Fnorm decreasing by nearly 45%, demonstrating that it does not meet the $O(3)$ equivariance requirements for tensor properties. The results on the elastic tensor dataset are shown in Table 5. Similar to the dielectric tensor dataset, the performance of the eComFormer method significantly declined on the augmented test set, with Fnorm dropping by nearly 29%. In contrast, GoeCTP’s performance remained unchanged, further demonstrating the effectiveness of the GoeCTP method. Additionally, runtime comparisons between GoeCTP and eComFormer on the test set showed that GoeCTP introduced no significant increase in runtime. This indicates that when integrated into single-value property prediction networks, GoeCTP incurs almost no additional computational cost, enhancing its practicality without compromising efficiency.
@@
Fnorm $\downarrow$ 3.23 4.71 3.23 3.23
EwT 25% $\uparrow$ 83.2% 69.7% 83.2% 83.2%
EwT 10% $\uparrow$ 56.8% 42.7% 56.8% 56.8%
EwT 5% $\uparrow$ 35.5% 22.5% 35.5% 35.5%
@@
Table 4: Ablation study for verifying the $O(3)$ equivariance with dielectric dataset.
Fnorm $\downarrow$ 107.11 138.45 107.11 107.11
EwT 25% $\uparrow$ 42.5% 25.9% 42.5% 42.5%
EwT 10% $\uparrow$ 15.3% 2.2% 15.3% 15.3%
EwT 5% $\uparrow$ 7.2% 0.2% 7.2% 7.2%
@@
Table 5: Ablation study for verifying the $O(3)$ equivariance with elastic dataset.
Efficiency. The results presented in Table 6 and Table 7 illustrate the running times for GoeCTP and baseline methods. On the dielectric dataset, GoeCTP completed the entire training process with only 38.2% of the time spent compared to GMTNet. On the elastic dataset, GoeCTP required less than 7.0% of the time spent by GMTNet to complete the entire training process. To achieve $O(3)$ equivariance for tensor properties, GMTNet’s network architecture relies on irreducible representations and tensor operations, which considerably reduce computational efficiency, especially for higher-order tensor property (elastic tensor) prediction tasks where the time cost of GMTNet increases sharply. In contrast, GoeCTP requires no special architectural design for tensor property prediction tasks; it predicts tensor properties of varying orders using only a multilayer perceptron (MLP) at the network’s output, ensuring both efficiency and scalability across different tensor orders.
@@
Time/batch (s) $\downarrow$ 0.052 0.104 0.126 0.048
@@
Time/batch (s) $\downarrow$ 0.226 0.348 $>$ 2.813 0.189
@@
6 Conclusion, limitations, and future works
In this work, we propose a novel $O(3)$-equivariant framework GoeCTP for fast crystal tensor prediction, which is a plug-and-play framework that can be integrated with any existing single-value property prediction network, enabling them to predict tensor properties with almost no additional computational cost. Based on predicting invariant tensor properties from standardized crystal positions and using an external module to ensure tensor equivariance, this approach has achieved state-of-the-art performance and the highest efficiency in different widely used crystal tensor property benchmark tests. The limitations of our current GoeCTP include (1) The performance of GoeCTP is inherently dependent on the performance of the Prediction module and, in theory, does not introduce additional performance enhancements. (2) Currently, GoeCTP is specifically tailored for the prediction of tensor properties in crystalline materials and has not yet been adapted for other material types. In future work, we intend to explore the following directions: for (1), we aim to improve performance by integrating prior knowledge related to the independent components of tensor properties across different crystal systems in Appendix A.2; for (2), we plan to extend the framework to other domains, such as 3D point cloud research, broadening its applicability.
@@ Appendix A Appendix
A.1 Details of eComFormer
@@
The eComFormer has demonstrated strong performance across a range of single-value property prediction tasks (Yan et al., 2024a). It is built on an SO(3)-equivariant crystal graph representation, where the interatomic distance vectors are employed to represent the edge features of the graph. The model converts this crystal graph into embeddings and utilizes a transformer architecture, incorporating both a node-wise transformer layer and a node-wise equivariant updating layer, to extract rich geometric information during the message-passing process.
@@
Figure 3: The detailed architectures of the node-wise transformer layer and node-wise equivariant updating layer, adapted from Yan et al. (2024a).
@@
where $\bm{\xi}_{K}$, $\bm{\xi}_{V}$ represent nonlinear transformations applied to key and value features, respectively, and the operators $\circ$ denote the Hadamard product, $\bm{\xi}_{msg}(\cdot)$ refers to the batch normalization layer, and $\sqrt{d_{\bm{q}_{ij}}}$ indicates the dimensionality of $\bm{q}_{ij}$. Then, node feature $\bm{f}_{i}$ is updated as follows,
@@
The node-wise equivariant updating layer is designed to effectively capture geometric features by incorporating node feature $\mathbf{a}_{i}$ and edge feature $||\mathbf{e}_{ji}||_{2}$ as inputs and stacking two tensor product (TP) layers (Geiger & Smidt, 2022). It uses node feature $\bm{f}_{i}^{l}$ and equivalent vector feature $\mathbf{e}_{ji}$ embedded by corresponding spherical harmonics $\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji})=c_{0}$,$\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji})=c_{1}*\frac{\mathbf{e}_{ji}}{||\mathbf{e}_{ji}||_{2}}\in\mathbb{R}^{3}$ and $\mathbf{Y}_{2}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{5}$ to represent the input features. Gathering rotational information from neighboring nodes to the central node $i$, the first TP layer is shown as

### v3

@@ 4.1 Experimental setup
Datasets. In this work, we evaluate the performance of GoeCTP on three key tensor property prediction tasks: the second-order dielectric tensor, the third-order piezoelectric tensor, and the fourth-order elastic tensor, respectively. Further details of the dataset and experimental setup can be found in Appendix A.10.
Baseline Methods. We selected several state-of-the-art methods in the field of crystal tensor property prediction, i.e. MEGNET [1, 31], EGTNN [53], and GMTNet [51], as baseline methods.
@@ 4.2 Experimental Results
Predicting Dielectric Tensors. The performance of various models in predicting the dielectric tensor is summarized in Table 1. While GoeCTP and GMTNet show close results in terms of Fnorm and EwT 25%, GoeCTP demonstrates higher values for EwT 10% and EwT 5%, with an improvement of 5% in EwT 5% compared to GMTNet. This indicates that GoeCTP delivers higher-quality predictions compared to GMTNet. Furthermore, as GoeCTP is designed to be a flexible framework, we also evaluated its combination with other models, such as iComFormer [50] and CrystalFormer [42]. Detailed results of these combinations are provided in Appendix A.11. In addition to these experimental results, we further investigate the impact of different canonical forms on tensor property prediction in crystalline materials. Detailed analyses can be found in Appendix A.16.
@@
Predicting Piezoelectric Tensors. The experimental results for the elastic tensor dataset are shown in Table 2. Similar to the experimental results on the dielectric tensors dataset, it achieves optimal results on all evaluation metrics. Further details on the results of GoeCTP’s combination with other models can be found in Appendix A.11.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Fnorm $\downarrow$ 0.504 0.515 0.450 0.431
EwT 25% $\uparrow$ 39.3% 36.3% 45.5% 46.9%
EwT 10% $\uparrow$ 27.1% 20.5% 38.1% 42.9%
EwT 5% $\uparrow$ 11.4% 13.0% 34.3% 39.7%
Table 2: Comparison of performance metrics between MEGNET, ETGNN, GMTNet, and GoeCTP on the piezoelectric dataset.
Predicting Elastic Tensors. Experimental results on the elastic tensor dataset are presented in Table 3. GoeCTP demonstrated outstanding performance in predicting higher-order, complex tensors, outperforming all baseline methods across every evaluation metric. Notably, it achieved the lowest Fnorm of 107.11 and improved all EwT metrics by an average of 6% compared to GMTNet, emphasizing its general applicability in predicting diverse tensor properties in materials science. Further details on the results of GoeCTP’s combination with other models can be found in Appendix A.11.
@@
Dielectric: Fnorm $\downarrow$ 3.23 4.71 3.23 3.23
@@
Piezoelectric: Fnorm $\downarrow$ 0.431 0.527 0.431 0.431
Total Time (s) $\downarrow$ 37.83 37.23 38.01 37.71
Elastic: Fnorm $\downarrow$ 107.11 138.45 107.11 107.11
@@
Table 4: Ablation study for verifying the $O(3)$ equivariance with dielectric, piezoelectric, and elastic dataset. After training GoeCTP, we extracted its Prediction module (i.e., eComFormer) for comparative testing on two different test sets
Verifying the $O(3)$ Equivariance. To evaluate the effectiveness of GoeCTP, we conducted experiments to verify the $O(3)$ equivariance of tensor properties. Specifically, after training GoeCTP, we extracted its Prediction module (i.e., eComFormer) for comparative testing on two different test sets (original test set and augmented test set). All crystals in the original test set were adjusted to the canonical form, while the augmented test set was generated by applying arbitrary $O(3)$ group transformations to all crystals in the original test set. The method for generating the corresponding orthogonal matrices is from Heiberger [10]. We then evaluated both the Prediction module (eComFormer) and GoeCTP on these two datasets and compared the performance metrics.
The results on the three tensor datasets are shown in Table 4, GoeCTP performed equally well on the augmented test set as on the original test set in all tensor datasets, indicating that it maintains strong $O(3)$ equivariance for tensor properties. In contrast, the performance of the eComFormer method significantly declined on the augmented test set, with Fnorm decreasing by nearly 45% in dielectric tensor dataset, nearly 22% in piezoelectric tensor dataset, and nearly 29% in elastic tensor dataset, demonstrating that it does not meet the $O(3)$ equivariance requirements for tensor properties.
@@
Time/batch (s) $\downarrow$ 0.043 0.085 0.103 0.039
@@
Additionally, runtime comparisons between GoeCTP and eComFormer on the test set revealed that GoeCTP did not result in a significant increase in runtime. This demonstrates that when integrated into single-value property prediction networks, GoeCTP incurs almost no additional computational cost, enhancing its practicality while maintaining efficiency.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Total Time (s) $\downarrow$ 943 1220 5771 938
Time/batch (s) $\downarrow$ 0.060 0.078 0.370 0.060
Table 6: Efficiency comparison on the piezoelectric dataset.
Efficiency. The results presented in Table 5, Table 6, and Table 7 illustrate the running times for GoeCTP and baseline methods. On the dielectric dataset, GoeCTP completed the entire training process with only 38.2% of the time spent compared to GMTNet. On the piezoelectric dataset, GoeCTP completed the entire training process with only 16.3% of the time spent compared to GMTNet. On the elastic dataset, GoeCTP required less than 7.0% of the time spent by GMTNet to complete the entire training process.
@@
Time/batch (s) $\downarrow$ 0.036 0.056 $>$ 0.452 0.030
@@
To achieve $O(3)$ equivariance for tensor properties, GMTNet’s network architecture relies on irreducible representations and tensor operations, which considerably reduce computational efficiency, especially for higher-order tensor property (elastic tensor) prediction tasks where the time cost of GMTNet increases sharply. In contrast, GoeCTP requires no special architectural design for tensor property prediction tasks; it predicts tensor properties of varying orders using only a MLP at the network’s output, ensuring both efficiency and scalability across different tensor orders.
5 Concluding Remarks
In this work, we revisit crystal tensor property prediction through the lens of canonicalization. Built upon the notion of canonicalization, we proposed a novel $O(3)$-equivariant framework GoeCTP for fast and accurate crystal tensor prediction. Benefiting from the canonicalization mechanism, GoeCTP serves as a plug-and-play module that can be integrated with any existing scalar property prediction network, enabling it to predict tensor properties with negligible additional computational cost. The proposed approach achieves state-of-the-art predictive performance and superior efficiency across a variety of widely used benchmark datasets for crystal tensor properties.
Except for the excel performance, our current GoeCTP has some limitations remain, which include (1) The performance of GoeCTP is inherently dependent on the performance of the Prediction module. (2) GoeCTP is specifically tailored for the prediction of tensor properties in crystalline materials and has not yet been adapted for other material types (See Appendix A.12 for more limitations details). In future study, we intend to improve performance by integrating prior knowledge related to the independent components of tensor properties across various crystal systems (see Appendix A.9). Further discussion on the utilization of space group constraints on tensor properties can be found in Appendix A.13.
@@ Appendix A Appendix
A.1 Related work
GNN-Based Methods. CGCNN is a pioneering GNN model specifically designed for handling crystal structures [47]. This model proposed to represent crystal structures as multi-edge crystal graphs. It was applied to predict various single-value properties such as formation energy and band gap. Since then, several GNN methods have been developed to improve upon CGCNN through exploring various network designs or leveraging prior knowledge [1, 26, 2, 3, 21]. These GNN methods are primarily designed for single-value property prediction and do not address the prediction of high-order tensor properties, such as dielectric or elastic tensors. Furthermore, they lack the ability to preserve the equivariance required for accurate high-order tensor property prediction. In contrast, recent studies attempted to ensure equivariance through specialized network architectures [28, 25, 11, 46, 34, 51, 53]. These approaches generally employ harmonic decomposition to achieve equivariance for tensor properties. Within these network architectures, many operations are required, such as tensor products and combining irreducible representations. These processes significantly increase computational costs, especially when handling higher-order data.
Transformer-Based Methods. Transformers, with their self-attention mechanism and parallel processing capabilities, are particularly well-suited for predicting crystal material properties. Matformer [49], one of the earliest Transformer-based networks used for crystal material property prediction, encoded crystal periodic patterns by using the geometric distances between the same atoms in neighboring unit cells. This addressed the issue in earlier GNN-based methods, including CGCNN, MEGNet, GATGNN, and others, which neglected the periodic patterns of infinite crystal structures. Subsequently, more advanced Transformer-based approaches were proposed. Some methods, such as ComFormer [50], CrystalFormer [44], and CrystalFormer [42], typically incorporate either enhanced graph construction techniques or physical priors, exhibiting impressive results in the prediction of single-value properties. Furthermore, DOSTransformer [20] is tailored for the density of states prediction, utilizing prompt-guided multi-modal transformer architecture to achieve super performance. Nevertheless, these models are not directly applicable for accurate high-order tensor property prediction due to their inability to capture the necessary equivariance.
Equivariant Learning. Recently, two general-purpose equivariant learning techniques for the arbitrary group have gained widespread attention: frame averaging [22, 36, 5] and canonicalization [27, 17, 6]. Frame averaging is developed based on group averaging and involves averaging over smaller frames instead of the entire group. Canonicalization, on the other hand, is inherently related to frame averaging and can be viewed as a more general form of it. [27]. In particular, Lin et al. [22] introduced minimal frame averaging, a method designed to achieve both efficient and exact equivalence in equivariant learning. This approach is primarily applicable to scenarios where the input and output spaces are identical, such as the n-body problem. It does not extend to cases involving different spaces, such as crystal tensor prediction. If this method were extended to tensor property prediction, it would represent a special case of Proposition 3.4, where both orbit canonicalization and rigid registration are achieved using QR decomposition. Within Proposition 3.4, however, different functions can be employed for orbit canonicalization and rigid registration. For instance, the Kabsch algorithm [18, 19] can be employed for rigid registration, while polar decomposition can be utilized for canonicalization.
To the best of our knowledge, no previous work has extended these equivariant learning techniques to the prediction of crystal tensor properties. Thus, our method represents the first extension of these concepts to the crystal tensor property prediction.
A.2 Details of Framework Architecture
Crystal Graph Construction. In GoeCTP, we use the crystal graph construction from Yan et al. [50, 51] to describe the structure and relationships within crystals. Specifically, we first convert the fractional coordinate system $(\mathbf{A},\mathbf{F},\mathbf{H})$ into the Cartesian coordinate system $(\mathbf{A},\mathbf{X},\mathbf{H})$ as introduced in Section 2. Then, we assume that the output crystal graph is represented as $\mathcal{G}(\mathcal{V},\mathcal{E})$, $\mathcal{V}$ denotes the set of nodes $v_{i}$ in the crystal graph, where each node $v_{i}$ contains atomic features $\mathbf{v}_{i}=(\bm{a}_{i},\hat{\bm{p}}_{i})$. $\mathcal{E}$ represents the set of edges denoted as $e_{ij}$, which are typically constructed based on the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$. When the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$ is less than a given radius $R$, i.e. $d_{ij}=||\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}||_{2}\leq R$, an edge $e_{ij}$ will be built with edge feature $\mathbf{e}_{ij}=\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}$. Here, $R$ is established based on the distance to the $k$-th nearest neighbor, and different values of $\bm{k}=[k_{1},k_{2},k_{3}]\in\mathbb{Z}^{3}$ represent different edges between nodes $v_{i}$ and $v_{j}$. Since there are no equivariance requirements for subsequent models, any other graph construction methods can be used to replace this part, such as Wang et al. [44], Yan et al. [50].
@@ A.2 Details of Framework Architecture
eComFormer. The eComFormer has demonstrated strong performance across a range of single-value property prediction tasks [50]. It is built on an SO(3)-equivariant crystal graph representation, where the interatomic distance vectors are employed to represent the edge features of the graph. The model converts this crystal graph into embeddings and utilizes a transformer architecture, incorporating both a node-wise transformer layer and a node-wise equivariant updating layer, to extract rich geometric information during the message-passing process.
@@
Figure 3: The detailed architectures of the node-wise transformer layer and node-wise equivariant updating layer, adapted from Yan et al. [50].
@@
where $\bm{\xi}_{K}$, $\bm{\xi}_{V}$ represent nonlinear transformations applied to key and value features, respectively. The operators $\circ$ denote the Hadamard product, $\operatorname{BN}(\cdot)$ refers to the batch normalization layer, and $\sqrt{d_{\bm{q}_{ij}}}$ indicates the dimensionality of $\bm{q}_{ij}$. Then, node feature $\bm{f}_{i}$ is updated as follows,
@@
The node-wise equivariant updating layer is designed to effectively capture geometric features by incorporating node feature $\mathbf{a}_{i}$ and edge feature $||\mathbf{e}_{ji}||_{2}$ as inputs and stacking two tensor product (TP) layers [7]. It uses node feature $\bm{f}_{i}^{l}$ and equivalent vector feature $\mathbf{e}_{ji}$ embedded by corresponding spherical harmonics $\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji})=c_{0}$,$\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji})=c_{1}*\frac{\mathbf{e}_{ji}}{||\mathbf{e}_{ji}||_{2}}\in\mathbb{R}^{3}$ and $\mathbf{Y}_{2}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{5}$ to represent the input features. Gathering rotational information from neighboring nodes to the central node $i$, the first TP layer is shown as